OpenStudy (anonymous):

dy/dx-y=e^x given that y(e) - 0

OpenStudy (epoweritheta):

\[y \prime -y = e^x\] \[-e^{-x}y \prime +e^{-x}y = -1\] \[(e^{-x}y)= -1\]

OpenStudy (epoweritheta):

edit last step : \[(e^{-x}y) \prime =-1\] \[e^{-x}y=-x +c\] finally

OpenStudy (solomonzelman):

In general, if you are given an equation in a form: \(\color{#000000}{ \displaystyle \frac{dy}{dx}+p'(x)=q(x) }\) Then, multiply the entire equation by the "integrating factor" (The integrating factor is: \(\color{red}{H(x)=e^{p(x)}}\)), \(\color{#000000}{ \displaystyle \color{red}{e^{p(x)}}\frac{dy}{dx}+\color{red}{e^{p(x)}}p'(x)=\color{red}{e^{p(x)}}q(x) }\) Note that, the left side is the derivative (by product rule) of \(\color{black}{y\cdot e^{p(x)}}\), so: \(\color{#000000}{ \displaystyle \frac{d}{dx}\left[y\cdot e^{p(x)}\right]=e^{p(x)}q(x) }\) Next integrate both sides. Suppose that, \(\color{#000000}{ \displaystyle \int e^{p(x)}q(x) ~dx=S(x)+C }\) And recall that, the anti-derivative of a derivative of a function is just that function itself. So, you are going to get. \(\color{#000000}{ \displaystyle \int \left(\frac{d}{dx}\left[y\cdot e^{p(x)}\right]\right)~dx=\int \left(e^{p(x)}q(x)\right)~dx }\) \(\color{#000000}{ \displaystyle y\cdot e^{p(x)}=S(x)+C }\) So, your explicit solution is: \(\color{#000000}{ \displaystyle y=e^{-p(x)}S(x)+Ce^{-p(x)} }\)

OpenStudy (solomonzelman):

The initial value problems don't differ much. You just plug in the given initial value to determine the C.